# On novel Hamiltonian description of the nonholonomic Suslov problem

### v1

@@ Abstract
We present some new Poisson bivectors that are invariants by the flow of the nonholonomic Suslov problem. Two rank four invariant Poisson bivectors have globally defined Casimir functions and, therefore, define cubic Poisson brackets of the five dimensional state space with standard symplectic leaves. For the Suslov gyrostat in the potential field we found rank two Poisson bivectors having only two globally defined Casimir functions and, therefore, we say about formal Hamiltonian description in these cases.
@@ 1 Introduction
with $I_{11}>0$ and $I_{22}>0$.
@@
where $k$ is a nonzero integer. In both cases exists third scalar solution $f_{3}(x)$ of the invariance equation (1.5) which can be found in [6].
@@ 1.1 Known invariants
$\mathcal{L}_{X}f=X^{1}\frac{\partial f}{\partial x_{1}}+\cdots+X^{6}\frac{\partial f}{\partial x_{6}}=0\,.$ (1.5)
We can solve this equation using method of undetermined coefficients and obtain energy and length of the Poisson vector $\gamma$ as a base in the space of solutions
@@
$\mbox{div}X=\sum_{k=1}^{5}\frac{\partial X_{k}}{x_{k}}=\frac{I_{23}}{I_{22}}\omega_{1}-\frac{I_{13}}{I_{11}}\omega_{2}\,.$ (1.7)
[…]
@@
$\Omega=D^{-1}(x)\,dx_{1}\wedge\cdots\wedge dx_{5}$ (1.10)
@@ 2 Rank four invariant Poisson bivectors
$[\![P,P]\!]=0$ (2.1)
which we represent using the Schouten-Nijenhuis bracket [[.,.]][\![.,.]\!][ [ . , . ] ] on multivector fields. Any bivector field can be regarded as a skew homomorphism and the rank of this field at a point $x_{0}$ is the rank of the induced linear mapping [23].
@@ 2.1 Cubic Poisson brackets
$(\mathcal{L}_{X}T)^{ij}=\sum_{k=1}^{6}\left(X^{k}\frac{\partial T^{ij}}{\partial x^{k}}-T^{kj}\frac{\partial X^{i}}{\partial x^{k}}-T^{ik}\frac{\partial X^{j}}{\partial x^{k}}\right)=0\,.$ (2.3)
We solve this equation using method of undetermined coefficients when we suppose that entries of multivector field $T(x)$ are inhomogeneous polynomials on variables $x_{1},\ldots,x_{5}$ of degree $N$.
At $N=1$ we have the following rank two invariant bivector
@@
Here $P_{1},P_{2}$ and $P_{3}$ are the Poisson bivectors which satisfy to the Jacobi condition (2.1). They are incompatible with each other and therefore $T_{3}$ is the skew symmetric multivector field of valence (2,0) which does not satisfy the Jacobi condition
$[\![T_{3},T_{3}]\!]\neq 0\,.$
This invariant bivector can be useful in the study of Hamiltonian systems on an almost-Poisson manifold, see [9, 4]. An almost-Poisson manifold is a manifold equipped with a skew-symmetric (2, 0)-tensor field that does not necessarily satisfy the Jacobi identity.
@@
The multiplication of two tensor invariants results in either a tensor invariant or zero. Therefore, we easy compute the following relations
@@
$X=P_{b}dH^{(1,2)}_{b}\,,\qquad H_{b}^{(1)}=\frac{\ln f_{1}}{2c_{5}}\,,\quad H_{b}^{(2)}=\frac{\ln f_{2}}{4c_{5}}\,.$
According to the basic result of Lie [15] if rank of the Poisson manifold is constant near point $x_{0}$, then there are coordinates $(q_{1},\ldots q_{n},p_{q_{1}},\ldots,p_{q_{k}},y_{1},\ldots,y_{s})$ near $x_{0}$ satisfying the canonical bracket relations
$\{q_{i},q_{j}\}=\{p_{i},p_{j}\}=\{q_{i},y_{j}\}=\{p_{i},y_{j}\}=\{y_{i},y_{j}\}=0\quad\mbox{and}\quad\{q_{i},p_{j}\}=\delta_{ij}\,.$ (2.10)
The coordinates $(q,p)$ are known as Darboux coordinates on the symplectic leaf $S$ which are only locally defined in general case.
It means that we can reduce Euler-Poisson equations (LABEL:ep-eq) to the Hamiltonian equations of motion
$\dot{q}_{i}=\frac{\partial H}{\partial p_{i}}\,,\qquad\dot{p}_{i}=-\frac{\partial H}{\partial q_{i}}\,,\qquad i=1,2,$
on the symplectic leaves of the Poisson bivectors $P_{a}$ or $P_{b}$ with $H=H_{a}$ and $H=H_{b}$, respectively. Then we can apply the standard theory of symplectic integrators to obtain numerical solutions of these equations [10].
3 Rank two invariant Poisson tensors
The classical heavy top problem admits multiple generalizations. To illustrate this point, consider the alternative of an arbitrary potential force field in lieu of a constant gravitational field. Another generalization is the so-called heavy gyrostat. The object is a rigid body that is subject to the forces of gravity and gyroscopic force.
Following [2, 17] we consider motion of the gyrostat in the potential field $U(\gamma_{1},\gamma_{2},\gamma_{3})$ defined by the following equations of motion
$\dot{\gamma}_{1}=-\omega_{2}\gamma_{3},\qquad\dot{\gamma}_{2}=\omega_{1}\gamma_{3},\qquad\dot{\gamma}_{3}=\omega_{2}\gamma_{1}-\omega_{1}\gamma_{2}$ (3.1)
$I_{11}\dot{\omega}_{1}=-(I_{13}\omega_{1}+I_{23}\omega_{2}+\Lambda_{3})\omega_{2}+\gamma_{2}\frac{\partial U}{\partial\gamma_{3}}-\gamma_{2}\frac{\partial U}{\partial\gamma_{2}}\,,$
$I_{22}\dot{\omega}_{2}=\phantom{-}(I_{13}\omega_{1}+I_{23}\omega_{2}+\Lambda_{3})\omega_{1}+\gamma_{3}\frac{\partial U}{\partial\gamma_{1}}-\gamma_{1}\frac{\partial U}{\partial\gamma_{3}}\,.$
Here $\Lambda_{3}$ is a nontrivial component of constant gyrostatic moment, see [17] for details.
3.1 Free gyrostat
If $U(\gamma)=0$, then system (3.1) has the same scalar invariants $f_{1,2}$ (1.6) and modified Darboux polynomial
$D_{\Lambda}(x)=D(x)+\lambda=I_{13}\omega_{1}+I_{23}\omega_{2}+\Lambda_{3}\,,$
see [17].
In the space of multivector fields of valency (2, 0) with inhomogeneous cubic entries the invariance equation (1.3) has the following generic rank four solution
$T_{3}(\Lambda)=(c_{1}f_{1}+c_{2}f_{2}+c_{3})P_{1}+c_{4}P_{2}(\Lambda)\,,\qquad c_{k}\in\mathbb{R}\,.$
Unlike (2.5), it depends only on four constants, and the Poisson bivector $P_{2}(\Lambda)$ is given by (2.6), where the polynomial $D(x)$ is replaced by $D_{\Lambda}(x)$.
The Jacobi condition (2.1) for $T_{3}$ is satisfied at $c_{4}=0$ or at $c_{1}=c_{2}=c_{3}=0$. In both cases we have rank two Poisson bivectors
$P_{c}=(c_{1}f_{1}+c_{2}f_{2}+c_{3})P_{1}\qquad\mbox{and}\qquad P_{d}=c_{4}P_{2}(\Lambda)$ (3.2)
@@ 3.1 Free gyrostat
$P_{c}df_{1}=0\,,\quad P_{c}df_{2}=0\qquad\mbox{and}\qquad P_{d}df_{1}=0\,,\quad P_{d}df_{2}=-2c_{4}f_{2}X\,.$ (3.3)
Obviously, we can not construct three independent globally defined invariant Casimir functions for these rank two bivectors using only two existing scalar invariants $f_{1,2}$.
In [17] the authors used method of the undetermined coefficients for the search of polynomial first integrals of the gyrostatic Suslov problem and obtained three scalar invariants:
  1. 1.
    if $I_{13}=0$ and $I_{11}=I_{22}+I_{23}^{2}/I_{22}$, then the additional first integral is
    $f_{3}=(I_{22}^{2}+I_{23}^{2})\omega_{1}\gamma_{1}+(\omega_{2}I_{22}^{2}-\Lambda_{3}I_{23})\gamma_{2}+I_{22}(\omega_{2}I_{23}+\Lambda_{3})\gamma_{3}$
  2. 2.
    if $I_{23}=0$ and $I_{22}=I_{11}+I_{13}^{2}/I_{11}$, then the additional first integral is
    $f_{3}=(I_{11}^{2}\omega_{1}-I_{13}^{2}\Lambda_{3})\gamma_{1}+(I_{11}^{2}+I_{13}^{2})\omega_{2}\gamma_{2}+I_{11}(\omega_{1}I_{13}+\Lambda_{3})\gamma_{3}$
  3. 3.
    if $I_{13}=I_{23}$ and $I_{22}=I_{11}$, then the additional first integral is
    $f_{3}=I_{11}\gamma_{1}\omega_{1}+I_{11}\gamma_{2}\omega_{2}+\Lambda_{3}\gamma_{3}\,.$
In all these case the system (3.1) is integrable in the Jacobi sense. We will consider only the first and third case, since the second case can easily be obtained from the first one by permuting the subscripts 1 and 2, see [2, 6].
One would expect that the existence of third independent scalar invariant would allow one to find the desired globally defined invariant Casimir functions. However, this does not happen in the these cases of integrability since an existence of the second invariant vector field.
Indeed, it is easy to prove that in all the cases of integrability the flow of $X$ preserves vector field $Y=P_{1}df_{3}$, i.e.
$\mathcal{L}_{X}Y=[X,Y]=0\,,$
where [.,.][.,.][ . , . ] is a Lie bracket. In the first case invariant vector field is equal to
$Y=P_{1}df_{3}=\left(\begin{array}[]{c}\omega_{2}\gamma_{3}I_{22}^{2}-\gamma_{2}I_{22}(\omega_{2}I_{23}+\Lambda_{3})-\Lambda_{3}\gamma_{3}I_{23}\\ -\omega_{1}\gamma_{3}I_{22}^{2}+\gamma_{1}I_{22}(\omega_{2}I_{23}+\Lambda_{3})-\omega_{1}\gamma_{3}I_{23}^{2}\\ \gamma_{2}(I_{22}^{2}+I_{23}^{2})\omega_{1}-\gamma_{1}(\omega_{2}I_{22}^{2}-\Lambda_{3}I_{23})\\ 0\\ 0\end{array}\right)\,,$
whereas in the third case it looks like
$Y=P_{1}df_{3}=\left(\begin{array}[]{c}I_{11}\omega_{2}\gamma_{3}-\Lambda_{3}\gamma_{2}\\ -I_{11}\omega_{1}\gamma_{3}+\Lambda_{3}\gamma_{1}\\ I_{11}(\omega_{1}\gamma_{2}-\omega_{2}\gamma_{1})\\ 0\\ 0\end{array}\right)\,.$
Thus, in all the cases we have
$P_{c}df_{3}=(c_{1}f_{1}+c_{2}f_{3}+c_{3})Y\qquad\mbox{and}\qquad P_{d}df_{3}=c_{4}f_{3}X$
in addition to (3.3). Formally the original vector field has a Hamiltonian form
$X=P_{d}H_{d}\,,\qquad H_{d}=-\frac{1}{2c_{4}}\ln f_{2}\,,$
with rank two Poisson bivector having only two globally defined Casimir functions
$P_{d}df_{1}=0\qquad P_{d}dC_{d}=0\,,\qquad C_{d}=\ln f_{2}+\ln f_{3}^{2}\,.$
It means that the Darboux coordinates on the corresponding symplectic leaves of $P_{d}$ are defined only locally.
3.2 Suslov system in the potential field
At $\Lambda_{3}=0$ equations (3.1) admit two scalar invariants
$f_{1}=I_{11}\omega_{1}^{2}+I_{22}\omega_{2}^{2}+2U(\gamma),\qquad f_{2}=\gamma_{1}^{2}+\gamma_{2}^{2}+\gamma_{3}^{2}.$ (3.4)
Divergency of vector field $X$ is equal to (1.7) but system (3.1) does not admit an invariant measure for general $U(\gamma)$ [17].
At the two special cases
$\alpha=0\,,\qquad U=\mu\,\ln\gamma_{3}\,,\qquad\mbox{and}\qquad\alpha\neq 0\,,\qquad U=\mu\gamma_{3}^{2\alpha}\,,\qquad\mu\in\mathbb{R}\,$
there exists tensor invariant in the space of multivector fields of valency (2,0)
$P_{e}=P_{2}+\alpha P_{3}+P_{\alpha}\,,$ (3.5)
@@ 3.2 Suslov system in the potential field
$P_{\alpha}=\left(\begin{array}[]{ccccc}0&0&0&-\frac{\gamma_{1}\gamma_{2}}{I_{11}}&\frac{\gamma_{1}^{2}}{I_{22}}\\ 0&0&0&\frac{-\gamma_{2}^{2}}{I_{11}}&\frac{\gamma_{1}\gamma_{2}}{I_{22}}\\ 0&0&0&-\frac{\gamma_{3}\gamma_{2}}{I_{11}}&\frac{\gamma_{3}\gamma_{1}}{I_{22}}\\ \frac{\gamma_{1}\gamma_{2}}{I_{11}}&\frac{\gamma_{2}^{2}}{I_{11}}&\frac{\gamma_{3}\gamma_{2}}{I_{11}}&0&\frac{\alpha(\omega_{1}\gamma_{1}I_{11}+\omega_{2}\gamma_{2}I_{22})}{I_{11}I_{22}}\\ -\frac{\gamma_{1}^{2}}{I_{22}}&-\frac{\gamma_{1}\gamma_{2}}{I_{22}}&-\frac{\gamma_{3}\gamma_{1}}{I_{22}}&-\frac{\alpha(\omega_{1}\gamma_{1}I_{11}+\omega_{2}\gamma_{2}I_{22})}{I_{11}I_{22}}&0\end{array}\right)\,\frac{\partial U}{\partial\gamma_{3}}\,.$
The multiplication of two tensor invariants results in either a tensor invariant or zero and we have
$P_{e}df_{1}=2f_{1}X\qquad\mbox{and}\qquad P_{e}df_{2}=2f_{2}X\,.$
The most important difference with bivectors $P_{a}$ (2.8) and $P_{b}$ (2.9) is that $P_{e}$ is a Poisson bivector of rank two. As noted earlier, we can not construct three independent globally defined Casimir functions from only two independent invariants $f_{1}$ and $f_{2}$. Thus, we have only a formal Hamiltonian description for the original vector field $X$
$X=P_{c}dH_{c}\,,\qquad\mbox{where}\qquad H_{c}=\frac{1}{2}f_{1,2}\,.$
We do not know how to describe the corresponding symplectic leaves and possible Darboux coordinates on them.
For the functions $U(\gamma_{1},\gamma_{2},\gamma_{3})$ of the special form we also have an invariant rank two Poisson bivectors and the corresponding formal Hamiltonian description, which we omit for brevity.

### v2

@@ Abstract
We present some new Poisson bivectors that are invariants by the flow of the nonholonomic Suslov problem. Two rank four invariant Poisson bivectors have globally defined Casimir functions and, therefore, define cubic Poisson brackets on the five dimensional state space with standard symplectic leaves. For the Suslov gyrostat in the potential field we found rank two Poisson bivectors having only two globally defined Casimir functions and, therefore, we say about formal Hamiltonian description in these cases.
@@ 1 Introduction
with $I_{11}>0$ and $I_{22}>0$. One more condition must be fulfilled for real bodies [6].
@@
where $k$ is a nonzero integer. In both cases exist three scalar solutions $f_{1}(x),f_{2}(x)$ and $f_{3}(x)$ of the invariance equation (1.5) which can be found in [10].
@@ 1.1 Known invariants
$\mathcal{L}_{X}f=X^{1}\frac{\partial f}{\partial x_{1}}+\cdots+X^{5}\frac{\partial f}{\partial x_{5}}=0\,.$ (1.5)
We can solve this equation using the method of undetermined coefficients and obtain the energy and length of the Poisson vector $\gamma$ as a base in the space of solutions
@@
$\mbox{div}X=\sum_{k=1}^{5}\frac{\partial X_{k}}{\partial x_{k}}=\frac{I_{23}}{I_{22}}\omega_{1}-\frac{I_{13}}{I_{11}}\omega_{2}\,.$ (1.7)
Substituting divXX into the definition of the Darboux polynomial $D(x)$ as a cofactor
@@
$\Omega=D(x)^{-1}\,dx_{1}\wedge\cdots\wedge dx_{5}$ (1.10)
@@ 2 Rank four invariant Poisson bivectors
[[P,P]]=0[\![P,P]\!]=0 (2.1)
which we represent using the Schouten-Nijenhuis bracket [[.,.]][\![.,.]\!] on multivector fields. Any bivector field can be regarded as a skew homomorphism and the rank of this field at a point $x_{0}$ is the rank of the induced linear mapping [27].
@@ 2.1 Cubic Poisson brackets
$(\mathcal{L}_{X}T)^{ij}=\sum_{k=1}^{5}\left(X^{k}\frac{\partial T^{ij}}{\partial x^{k}}-T^{kj}\frac{\partial X^{i}}{\partial x^{k}}-T^{ik}\frac{\partial X^{j}}{\partial x^{k}}\right)=0\,.$ (2.3)
We solve this equation using the method of undetermined coefficients when we suppose that the entries of the multivector field $T(x)$ are inhomogeneous polynomials on variables $x_{1},\ldots,x_{5}$ of degree $N$.
At $N=1$ we have the following known rank two invariant bivector
@@
Here $P_{1},P_{2}$ and $P_{3}$ are the Poisson bivectors which satisfy to the Jacobi condition (2.1). They are incompatible with each other and, therefore, $T_{3}$ is the symmetric skew multivector field of valence (2,0) which does not satisfy the Jacobi condition
$T_{3},T_{3}]\!]\neq 0\,.$
Nevertheless it is geometric structure which is invariant with respect to the Suslov model flow.
@@
A tensor operation between invariant tensor fields results in either a tensor invariant or zero. Therefore, we easy compute the following relations
@@
$X=P_{b}dH^{(1,2)}_{b}\,,\qquad H_{b}^{(1)}=\frac{\ln f_{1}}{2c_{5}}\,,\quad H_{b}^{(2)}=-\frac{\ln f_{2}}{4c_{5}}\,.$
3 Special case of the Euler-Poisson equations
The Euler equation describes the evolution of the angular velocity vector $\omega=(\omega_{1},\omega_{2},\omega_{3})$
$I\dot{\omega}+\omega\times I\omega=\mu r\times\gamma\,$,
Here $r=(r_{1},r_{2},r_{3})$ and $\gamma=(\gamma_{1},\gamma_{2},\gamma_{3})$ are radius vector of the center of mass and unit Poisson vector in the system of principal axes, $I=\mbox{diag}(I_{1},I_{2},I_{3})$ is a tensor of inertia, $\mu=mg$ is a weight of the body and $\times$ means a vector product, see all the details in textbook [3].
The Poisson equation describes evolution of the Poisson vector $\gamma$
$\dot{\gamma}=\gamma\times\omega\,$,
For a free rigid body $\mu=0$ and the Euler equation becomes independent of the Poisson equation. The Euler-Poisson equations for the Suslov problem (LABEL:ep-eq) are another example of equations that can be separated into two parts.
The third example is the Euler-Poisson equations
$I\dot{\omega}=\omega\times A\omega\qquad\mbox{and}\qquad\dot{\gamma}=\gamma\times B\omega\,$, (3.1)
where $A,B$ are two constant $3\times 3$ matrices. The system of equations (3.1) describes the partial motion of a rigid body with a fluid-filled cavity, see §3 of Chapter 3 in [3]. This is the case when the intensity of the fluid vortex in the cavity is small compared to the kinetic moment, or vice versa. These equations also arise in other physical applications, see [4, 5, 7] and references within.
The first equation in (3.1) is independent of the second equation which has zero divergency
$\sum_{i=1}^{3}\frac{\partial\left(\gamma\times BM\right)_{i}}{\partial\gamma_{i}}=0\,$,
so the phase flow of $X$ preserves the corresponding trivector $\mathcal{E}_{\gamma}$, vector field $Y$ and scalar invariant $F_{1}$ defined by
$\mathcal{E}_{\gamma}=\dfrac{\partial}{\partial\gamma_{1}}\wedge\dfrac{\partial}{\partial\gamma_{2}}\wedge\dfrac{\partial}{\partial\gamma_{3}}\,,\qquad Y=\sum_{i=1}^{3}\gamma_{i}\dfrac{\partial}{\partial\gamma_{i}}\,,\qquad F_{1}=(\gamma,\gamma)\,.$
If $A$ is a symmetric matrix, $A=A^{T}$, the divergency of the vector field $X$ (3.1) is equal to zero
$\mbox{div}\,X=\sum_{i=1}^{3}\frac{\partial\left(M\times AM\right)_{i}}{\partial M_{i}}+\sum_{i=1}^{3}\frac{\partial\left(\gamma\times BM\right)_{i}}{\partial\gamma_{i}}\,$,
and, therefore, the phase flow of $X$ preserves the volume form
$\Omega=d\gamma_{1}\wedge d\gamma_{2}\wedge d\gamma_{3}\wedge dM_{1}\wedge dM_{2}\wedge dM_{3}\,.$ (3.2)
In this case, the first equation in (3.1) coincides with the Euler equation describing free rigid body motion
$\dot{M}=M\times\omega\,,\qquad\mbox{where}\qquad\omega=AM\,$,
having two scalar invariants
$F_{2}=(M,M)\,,\qquad F_{3}=\frac{1}{2}(M,AM)\,.$
Substituting the well-known solutions $M(t)$ of the Euler equation into the second equation in (3.1) we non-autonomous Carathéodory differential equations of the form
$\dot{\gamma}=\gamma\times BM(t)\,.$
In the mathematical theory of control such systems are known as linear switched systems of differential equations.
According to the Euler–Jacobi–Lie theorem [17] if the system of ordinary differential equations
$\dot{x}=X(x_{1},\ldots,x_{n})$
on the $n$-dimensional smooth phase space with coordinates $x=(x_{1},\ldots,x_{n})$ possesses
  1. 1.
    $k$ functionally independent first integrals $F_{1},\ldots,F_{k}$;
  2. 2.
    m=nk2m=n-k-2 independent invariant symmetry fields $Y_{k+1},\ldots,Y_{n-2}$ which together with the field $Y_{n-1}=X$ generate a nilpotent Lie algebra of vector fields;
  3. 3.
    an invariant volume $n$-form $\Omega$;
    $\mathcal{L}_{X}\Omega=0\,.$
@@ 3 Special case of the Euler-Poisson equations
$\mathcal{L}_{Y_{i}}F_{j}=0,\qquad\mathcal{L}_{Y_{i}}\Omega=0\,$,
then it is integrated by quadratures.
In our case, we have $k=3$ independent first integrals $F_{1}$, $F_{2}$ and $F_{3}$, m=632=1m=6-3-2=1 vector field $Y$ commuting with $X$ and invariant volume form $\Omega$ (3.2). However, the vector field $X$ (3.1) does not satisfy the conditions of the Euler–Jacobi–Lie theorem since
$\mathcal{L}_{Y}F_{1}=2F_{1}\,.$
So, in the generic case, we have a non-integrable system.
3.1 Formal and informal Hamiltonization
We have the following tensor invariants of $X$ (3.1) :
  • three scalar invariants
    $F_{1}=(\gamma,\gamma)\,,\qquad F_{2}=(M,M)\,,\qquad F_{3}=\frac{1}{2}(M,AM)\,;$ (3.3)
  • three invariant multivectors
    $Y=\sum_{i=1}^{3}\gamma_{i}\dfrac{\partial}{\partial\gamma_{i}}\,,\qquad\mathcal{E}_{\gamma}=\dfrac{\partial}{\partial\gamma_{1}}\wedge\dfrac{\partial}{\partial\gamma_{2}}\wedge\dfrac{\partial}{\partial\gamma_{3}}\,$, (3.4)
    and
    $\mathcal{E}=\dfrac{\partial}{\partial\gamma_{1}}\wedge\dfrac{\partial}{\partial\gamma_{2}}\wedge\dfrac{\partial}{\partial\gamma_{3}}\wedge\dfrac{\partial}{\partial M_{1}}\wedge\dfrac{\partial}{\partial M_{2}}\wedge\dfrac{\partial}{\partial M_{3}}\,.$
The Lie derivatives $\mathcal{L}_{X}$ of all these invariants are equal to zero. We will use these invariants to construct Hamilton representation of $X$
$X=PdH$
where $H$ is a Hamilton function and $P$ is a Poisson bivector.
Another main ingredient is the semi invariant or Darboux invariant $Z$ of the homogeneous vector field $X$
$\mathcal{L}_{X}Z=X\,$,
where vector field
$Z=\sum_{i=1}^{3}\gamma_{i}\dfrac{\partial}{\partial\gamma_{i}}+\sum_{i=1}^{3}M_{i}\dfrac{\partial}{\partial M_{i}}$ (3.5)
is the well-known Euler vector field whose flow is the linear multiplication action. Since the first integrals $F_{k}$ are homogeneous functions we have
$\mathcal{L}_{Z}F_{k}=2F_{k}\,,\qquad k=1,2,3,$
according to the Euler theorem on homogeneous functions. Moreover, the invariant rank-two bivector
$P_{e}=X\wedge Z\,,\qquad\mathcal{L}_{X}P=0\,$,
satisfies to the Jacobi condition
$P_{e},P_{e}]\!]=0\,$,
according to the same Euler theorem.
The invariant bivector $P$ does not have four independent Casimir functions since we have only three single-valued scalar invariants $F_{1},F_{2}$ and $F_{3}$, see discussion in [1]. So, we can say that this rank-two bivector yields formal Hamiltonization of $X$.
It is simple to confirm that the rank four bivector
$P=(c_{1}F_{1}+c_{2}F_{2}+c_{3}F_{3})P_{\gamma}+c_{4}P_{c}+eP_{e}\,,\qquad c_{k},e\in\mathbb{R}\,$, (3.6)
@@ 3.1 Formal and informal Hamiltonization
$P_{\gamma}=\mathcal{E}_{\gamma}dF_{1}\,,\qquad P_{c}=X\wedge Y\qquad\mbox{and}\qquad P_{e}=X\wedge Z\,.$
is the invariant bivector
$\mathcal{L}_{X}P=0\,$,
which satisfies the Jacobi condition
[[P,P]]=0,[\![P,P]\!]=0\,,
and the following equations hold
$PdF_{1}=2(c_{4}+e)F_{1}X\,,\qquad PdF_{2}=2eF_{2}X\,,\qquad PdF_{3}=2eF_{3}X\,.$ (3.7)
From (3.7) follows that we can define two independent Casimir functions
$C_{1}=\left(1+\frac{c_{4}}{e}\right)\ln F_{2}-\ln F_{1}\,,\qquad C_{2}=\left(1+\frac{c_{4}}{e}\right)\ln F_{3}-\ln F_{1}\,.$
This means that $P_{D}C_{1},2=0$, and the vector field $X$ (3.1) can be rewritten in Hamiltonian form
$X=PdH\,,\qquad H=\frac{1}{e}\ln F_{2}\,.$
Here $P$ is a rank four Poisson bivector with two well-defined Casimir functions $C_{1,2}$ and $F_{2}$ is a mechanical energy. We can therefore talk about the informal Hamiltonisation of the vector field $X$.
If $A$ is an arbitrary matrix $A\neq A^{T}$, the divergence of the vector field $X$ (3.1) does not equal zero.
$\mbox{div}\,X=(a_{2,3}-a_{3,2})M_{1}+(a_{3,1}-a_{1,3})M_{2}+(a_{1,2}-a_{2,})M_{3}$
and, therefore, the phase flow of $X$ does not preserve the volume form $\Omega$ (3.2) and function
$F_{3}=\frac{1}{2}(M,AM)\,.$
The rank four invariant Poisson bivector (3.6) now depends on four parameters $c_{k}$
$P=(c_{1}F_{1}+c_{2}F_{2})P_{\gamma}+c_{3}P_{c}+eP_{e}\,,\qquad c_{k},e\in\mathbb{R}\,.$ (3.8)
But we cannot now construct two independent single-valued Casimir functions $C_{1,2}$ and the Hamiltonian $H$ from only two scalar invariants $F_{1,2}$ of the phase flow. So, in this case, we can only say about the formal Hamiltonization of the vector field $X$.